OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)].

OpenStudy (bahrom7893):

g(-6) = 1/(2*(-6)-2) = 1/(-12-2) = -1/14

OpenStudy (bahrom7893):

f [g(-6)] = f [ -1/14] = (-1/14) - 4

OpenStudy (anonymous):

-4 isnt an option.

OpenStudy (anonymous):

9 3 −9 −3

OpenStudy (bahrom7893):

Well did I say 4 is the answer? whatever -1/14 - 4 is.. wait

OpenStudy (bahrom7893):

By the way is it: